# No Dimensional Sampling Coresets for Classification

### v1

@@ 4.3 Coreset for svm Function
Proof  For a fixed $\alpha>0$, set $t=\alpha z$, $K=\alpha^{2}k$, and
@@
One can simply check that $\max_{t\geq 1}f_{K}(t)=1+2K$. Also, if we assume that $K\geq 1$, then, by simple calculation, we can derive that $\max_{0\leq t\leq 1}f_{K}(t)=f_{K}(1)=1+2K$. So, using Lemma 19, we obtain
@@
Replacing $K$ with $\alpha^{2}z$ concludes the result.
Lemma 35
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$. Then $s(y)=6+4kA^{2}$ is an upper sensitivity function for $(\mathcal{X},\ P,\ \mathcal{L}^{H}_{\operatorname{{\rm svm}},k})$ and $(\mathcal{X},\ P,\ \bar{\mathcal{L}}^{H}_{\operatorname{{\rm svm}},k})$.
Proof  Since $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$, our computation of $s(y)$ is limited to $\|y\|_{H}\leq A$. For $w\in\mathcal{W}=\{w\in\mathcal{X}\colon\|w\|_{H}\leq\frac{1}{A}\}$, Lemma 34 indicates that Lemma 20 is applicable with $B_{1}=A,B_{2}=\frac{1}{A},M=\phi(-1)=2$, and $\beta_{\operatorname{{\rm svm}}}(\alpha)=1+2\max(1,\alpha^{2}k).$ Using Lemma 21 with $\gamma(\|x\|_{H})=\frac{\operatorname{{\rm svm}}(0)}{M\beta_{\operatorname{{\rm svm}}}(\|x\|_{H})}=\frac{1}{2\beta_{\operatorname{{\rm svm}}}(\|x\|_{H})}$, we conclude that
@@
So, $s(y)=6+4kA^{2}$ is an upper sensitivity function for $(\mathcal{X},\ P,\ \mathcal{L}^{H}_{\operatorname{{\rm svm}},k})$.
Theorem 36
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$. For $S=6+4kA^{2}$, $C=(4A+2)\max(2Ak,1)+8Ak+1$, and $m=\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$, any iid sample $x_{1},\ldots,x_{m}$ from $\mathcal{X}$ with weights $u_{i}=\frac{1}{m}$ provides an $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{X},\mathcal{L}^{H}_{\operatorname{{\rm svm}},k})$ (respectively for $(\mathcal{X},P,\mathcal{X},\bar{\ell}^{H}_{\operatorname{{\rm svm}},k})$) with probability at least $1-\delta$.
Proof  Lemma 39 indicates that $s(y)=6+4kA^{2}$ is an upper sensitivity function. As it is a constant function, $s$-sensitivity sampling is the same as sampling according to $P$. Notice that $\operatorname{{\rm svm}}$ is convex and non-increasing. For $E_{1}=A$, using Equation 6, we have
@@
This implies that Definition 8 is satisfied for $\mathcal{L}^{H}_{\operatorname{{\rm svm}},k}$ and thus Theorem 10 concludes the statement for $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+\log\frac{1}{\delta}\right)$.
@@
Corollary 37 (Theorem 12, Restated)
Let $\boldsymbol{K}:\mathbb{R}^{d}\times\mathbb{R}^{d}\longrightarrow(0,1]$ be a reproducing kernel, i.e., a kernel associated with an RKHS, and $P$ be a probability measure over $\mathbb{R}^{d}$. For $\varepsilon,\delta\in(0,1)$, there exists a universal constant $C$ (independent of $d$ and $\boldsymbol{K}$) such that if $m\geq\frac{C}{\varepsilon^{2}}\log\frac{1}{\delta}$, then, with probability at least $1-\delta$, for any random sample $X=\{x_{1},\ldots,x_{m}\}$ based on $P$, we have
[…]
Proof  Note that $\boldsymbol{K}$ and $\phi(t)=\max(0,1-t)$ satisfy Definition 8 with $E_{1}=1$, $E_{2}=\frac{1}{2}$, and
$\ell^{H}_{\phi,k}(x,w)=\phi(\boldsymbol{K}(x,w))+\frac{1}{k}\|w\|_{H}^{2}=1-\boldsymbol{K}(x,w)+\frac{1}{k}\|w\|_{H}^{2}\leq 1-\boldsymbol{K}(x,w)+\frac{1}{k}.$
@@
$\left|\int_{x\in\mathbb{R}^{d}}\boldsymbol{K}(x,w){\rm d}P(x)-\frac{1}{m}\sum_{i=1}^{m}\boldsymbol{K}(x_{i},w)\right|$ $=\left|\int_{x\in\mathbb{R}^{d}}\ell^{H}_{\phi,1}(x,w){\rm d}P(x)-\frac{1}{m}\sum_{i=1}^{m}\ell^{H}_{\phi,1}(x_{i},w)\right|$
(by Theorem 36) $\leq\frac{\varepsilon}{2}\int_{x\in\mathbb{R}^{d}}\ell^{H}_{\phi,1}(x,w){\rm d}P(x)$
$\leq\frac{\varepsilon}{2}\int_{x\in\mathbb{R}^{d}}(2-\boldsymbol{K}(x,w)){\rm d}P(x)$
(since $\boldsymbol{K}(x,w)\leq 1$) $\leq\varepsilon.$
Theorem 38
@@
Proof  Given that Lemma 39 asserts the constant function $s(x)=3+2kA^{2}+\sqrt{k}A$ as an upper sensitivity function for $\mathcal{L}_{\operatorname{{\rm svm}},k}$, we deduce $\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm svm}},k}},\succ)=\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm svm}},k},\succ)$. Notice $\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm svm}},k}},\succ)=\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm svm}},k},\succ)$ since each function in $\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm svm}},k}}$ is a function in $\mathcal{L}_{\operatorname{{\rm svm}},k}$ scaled by a positive constant. For an $f_{w}\in\mathcal{L}_{\operatorname{{\rm svm}},k}$ and $r\geq 0$,
@@
which concludes that $\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm svm}},k},\succ)$ only includes half-spaces and the whole space $\mathbb{R}^{d}$. Therefore, by Radon’s theorem, ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm svm}},k},\succ))\leq d+1$ (for a proof, see Lemma 10.3.1 in  (Matousek, 2002)). Using Theorem 3, we also have the statement of Theorem 28 for $m\geq O(\frac{S}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right))$, completing the proof. A similar argument establishes the result for $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm svm}},k})$.
@@
Lemma 39
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}_{x\sim p}\|x\|_{H}\leq E_{1}^{2}$. Then $s(x)=(6+8kE_{1}^{2})\left(2+\frac{\|y\|_{H}}{E_{1}}\right)$ is an upper sensitivity function for $(\mathcal{X},P,\mathcal{L}^{H}_{\operatorname{{\rm svm}},k})$ and $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm svm}},k})$ with total sensitivity at most $18+24kE_{1}^{2}$.
Proof  As $\operatorname*{\mathbb{E}}_{x\sim P}(\|x\|_{H}^{2})\leq E^{2}_{1}$, Markov’s inequality implies $\operatorname*{\mathbb{P}}_{x\sim P}(\|x\|_{H}^{2}\geq 2E^{2}_{1})\leq\frac{1}{2}$. For an arbitrary $y\in\mathcal{X}$, define $\mathcal{X}_{y}=\left\{x\in\mathcal{X}\colon\|x\|_{H}\leq\sqrt{2}\max(E_{1},\|y\|_{H})\right\}$. Additionally, define $\mathcal{W}=\{w\in\mathcal{X}\colon\|w\|_{H}\leq\frac{1}{\sqrt{2}E_{1}}\}$. Lemma 26 indicates that Lemma 20 is applicable with $B_{1}=\sqrt{2}\max(E_{1},\|y\|_{H})$, $B_{2}=\frac{1}{\sqrt{2}E_{1}}$, $\operatorname{{\rm svm}}(-B_{1}B_{2})=1+\max\left(1,\frac{\|y\|_{H}}{E_{1}}\right)\leq 2+\frac{\|y\|_{H}}{E_{1}}=M_{y}$, and
@@
(Lemma 20) $\leq\sup_{\|w\|_{H}\leq B_{2}}\frac{1}{\int_{x\in\mathcal{X}}\frac{\operatorname{{\rm svm}}(0)}{M_{y}\beta(\|x\|_{H})}{\rm d}P(x)}$
@@
(*(*$\quad$( * ) $\leq(6+8kE_{1}^{2})\left(2+\frac{\|y\|_{H}}{E_{1}}\right).$
To see (***), note that
@@
With a similar proof, we can restate the following versions of Theorems 31 and 33 replacing $\operatorname{{\rm logistic}}$ by $\operatorname{{\rm svm}}$.
Theorem 40
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E_{1}$,
@@
4.4 Coreset for ReLU Function
@@
For $\text{ReLU}(t)=\max(0,t)$, we derive our results from those in previous section using a simple trick. Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E^{2}_{1}$ or $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$. It is easy to verify that $\boldsymbol{K}^{\prime}(\cdot)=1+\boldsymbol{K}(\cdot)$ is also a reproducing kernel associated with $H^{\prime}$ with $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H^{\prime}}^{2}\leq(1+E_{1})^{2}$ or $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H^{\prime}}\geq 1+A\right\}\right)=0$. As $\operatorname{{\rm svm}}(1+t)=\max(0,-t)$, if we plug $\boldsymbol{K}^{\prime}$ in the results given in the previous section, we can replicate them for $\phi(t)=\max(0,-t)$ by replacing $E_{1}$ and $A$ with $(1+E_{1})$ and $(1+A)$. While $\phi$ is not precisely the ReLU function, for the case that the kernel is the standard linear kernel in Euclidean space,
$\ell_{\phi,k}(x,w)=\max(0,-\langle x,w\rangle)+\frac{1}{k}\|w\|_{2}^{2}=\max(0,\langle x,-w\rangle)+\frac{1}{k}\|w\|_{2}^{2}=\ell_{\text{ReLU},k}(x,-w).$

### v2

@@ D.3 Coreset for svm Function
Proof.
For a fixed $\alpha>0$, set $t=\alpha z$, $K=\alpha^{2}k$, and
@@
One can simply check that $\max_{t\geq 1}f_{K}(t)=1+2K$. Also, if we assume that $K\geq 1$, then, by simple calculation, we can derive that $\max_{0\leq t\leq 1}f_{K}(t)=f_{K}(1)=1+2K$. So, using Lemma D.1, we obtain
@@
Replacing $K$ with $\alpha^{2}z$ concludes the result. ∎
Lemma D.17.
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$. Then $s(y)=6+4kA^{2}$ is an upper sensitivity function for $(\mathcal{X},\ P,\ \mathcal{L}^{H}_{\operatorname{{\rm svm}},k})$ and $(\mathcal{X},\ P,\ \bar{\mathcal{L}}^{H}_{\operatorname{{\rm svm}},k})$.
Proof.
Since $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$, our computation of $s(y)$ is limited to $\|y\|_{H}\leq A$. For $w\in\mathcal{W}=\{w\in\mathcal{X}\colon\|w\|_{H}\leq\frac{1}{A}\}$, Lemma D.16 indicates that Lemma D.2 is applicable with $B_{1}=A,B_{2}=\frac{1}{A},M=\phi(-1)=2$, and $\beta_{\operatorname{{\rm svm}}}(\alpha)=1+2\max(1,\alpha^{2}k).$ Using Lemma D.3 with $\gamma(\|x\|_{H})=\frac{\operatorname{{\rm svm}}(0)}{M\beta_{\operatorname{{\rm svm}}}(\|x\|_{H})}=\frac{1}{2\beta_{\operatorname{{\rm svm}}}(\|x\|_{H})}$, we conclude that
@@
So, $s(y)=6+4kA^{2}$ is an upper sensitivity function for $(\mathcal{X},\ P,\ \mathcal{L}^{H}_{\operatorname{{\rm svm}},k})$. ∎
Theorem D.18.
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$. For $S=6+4kA^{2}$, $C=(4A+2)\max(2Ak,1)+8Ak+1$, and $m=\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$, any iid sample $x_{1},\ldots,x_{m}$ from $\mathcal{X}$ with weights $u_{i}=\frac{1}{m}$ provides an $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{X},\mathcal{L}^{H}_{\operatorname{{\rm svm}},k})$ (respectively for $(\mathcal{X},P,\mathcal{X},\bar{\ell}^{H}_{\operatorname{{\rm svm}},k})$) with probability at least $1-\delta$.
Proof.
Lemma D.21 indicates that $s(y)=6+4kA^{2}$ is an upper sensitivity function. As it is a constant function, $s$-sensitivity sampling is the same as sampling according to $P$. Notice that $\operatorname{{\rm svm}}$ is convex and non-increasing. For $E_{1}=A$, using Equation 6, we have
@@
This implies that Definition A.1 is satisfied for $\mathcal{L}^{H}_{\operatorname{{\rm svm}},k}$ and thus Theorem A.3 concludes the statement for $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+\log\frac{1}{\delta}\right)$. ∎
@@
Corollary D.19 (Theorem A.5, Restated).
Let $\mathsf{K}:\mathbb{R}^{d}\times\mathbb{R}^{d}\longrightarrow(0,1]$ be a reproducing kernel, i.e., a kernel associated with an RKHS, and $P$ be a probability measure over $\mathbb{R}^{d}$. For $\varepsilon,\delta\in(0,1)$, there exists a universal constant $C$ (independent of $d$ and $\mathsf{K}$) such that if $m\geq\frac{C}{\varepsilon^{2}}\log\frac{1}{\delta}$, then, with probability at least $1-\delta$, for any random sample $X=\{x_{1},\ldots,x_{m}\}$ based on $P$, we have
$\sup_{w\in\mathbb{R}^{d}}|{\rm KDE}_{P}(w)-{\rm KDE}_{X}(w)|\leq\varepsilon.$
Proof.
Note that $\mathsf{K}$ and $\phi(t)=\max(0,1-t)$ satisfy Definition A.1 with $E_{1}=1$, $E_{2}=\frac{1}{2}$, and
$\ell^{H}_{\phi,k}(x,w)=\phi(\mathsf{K}(x,w))+\frac{1}{k}\|w\|_{H}^{2}=1-\mathsf{K}(x,w)+\frac{1}{k}\|w\|_{H}^{2}\leq 1-\mathsf{K}(x,w)+\frac{1}{k}.$
@@
$\left|\int_{x\in\mathbb{R}^{d}}\mathsf{K}(x,w){\rm d}P(x)-\frac{1}{m}\sum_{i=1}^{m}\mathsf{K}(x_{i},w)\right|$ $=\left|\int_{x\in\mathbb{R}^{d}}\ell^{H}_{\phi,1}(x,w){\rm d}P(x)-\frac{1}{m}\sum_{i=1}^{m}\ell^{H}_{\phi,1}(x_{i},w)\right|$
(by Theorem D.18) $\leq\frac{\varepsilon}{2}\int_{x\in\mathbb{R}^{d}}\ell^{H}_{\phi,1}(x,w){\rm d}P(x)$
$\leq\frac{\varepsilon}{2}\int_{x\in\mathbb{R}^{d}}(2-\mathsf{K}(x,w)){\rm d}P(x)$
(since $\mathsf{K}(x,w)\leq 1$) $\leq\varepsilon.$
Theorem D.20.
@@
Proof.
Given that Lemma D.21 asserts the constant function $s(x)=3+2kA^{2}+\sqrt{k}A$ as an upper sensitivity function for $\mathcal{L}_{\operatorname{{\rm svm}},k}$, we deduce $\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm svm}},k}},\succ)=\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm svm}},k},\succ)$. Notice $\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm svm}},k}},\succ)=\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm svm}},k},\succ)$ since each function in $\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm svm}},k}}$ is a function in $\mathcal{L}_{\operatorname{{\rm svm}},k}$ scaled by a positive constant. For an $f_{w}\in\mathcal{L}_{\operatorname{{\rm svm}},k}$ and $r\geq 0$,
@@
which concludes that $\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm svm}},k},\succ)$ only includes half-spaces and the whole space $\mathbb{R}^{d}$. Therefore, by Radon’s theorem, ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm svm}},k},\succ))\leq d+1$ (for a proof, see Lemma 10.3.1 in  (Matousek, 2002)). Using Theorem 1.2, we also have the statement of Theorem D.10 for $m\geq O(\frac{S}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right))$, completing the proof. A similar argument establishes the result for $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm svm}},k})$. ∎
@@
Lemma D.21.
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}_{x\sim p}\|x\|_{H}\leq E_{1}^{2}$. Then $s(x)=(6+8kE_{1}^{2})\left(2+\frac{\|y\|_{H}}{E_{1}}\right)$ is an upper sensitivity function for $(\mathcal{X},P,\mathcal{L}^{H}_{\operatorname{{\rm svm}},k})$ and $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm svm}},k})$ with total sensitivity at most $18+24kE_{1}^{2}$.
Proof.
As $\operatorname*{\mathbb{E}}_{x\sim P}(\|x\|_{H}^{2})\leq E^{2}_{1}$, Markov’s inequality implies $\operatorname*{\mathbb{P}}_{x\sim P}(\|x\|_{H}^{2}\geq 2E^{2}_{1})\leq\frac{1}{2}$. For an arbitrary $y\in\mathcal{X}$, define $\mathcal{X}_{y}=\left\{x\in\mathcal{X}\colon\|x\|_{H}\leq\sqrt{2}\max(E_{1},\|y\|_{H})\right\}$. Additionally, define $\mathcal{W}=\{w\in\mathcal{X}\colon\|w\|_{H}\leq\frac{1}{\sqrt{2}E_{1}}\}$. Lemma D.8 indicates that Lemma D.2 is applicable with $B_{1}=\sqrt{2}\max(E_{1},\|y\|_{H})$, $B_{2}=\frac{1}{\sqrt{2}E_{1}}$, $\operatorname{{\rm svm}}(-B_{1}B_{2})=1+\max\left(1,\frac{\|y\|_{H}}{E_{1}}\right)\leq 2+\frac{\|y\|_{H}}{E_{1}}=M_{y}$, and
@@
(Lemma D.2) $\leq\sup_{\|w\|_{H}\leq B_{2}}\frac{1}{\int_{x\in\mathcal{X}}\frac{\operatorname{{\rm svm}}(0)}{M_{y}\beta(\|x\|_{H})}{\rm d}P(x)}$
@@
((*$\quad$( ∗ ) $\leq(6+8kE_{1}^{2})\left(2+\frac{\|y\|_{H}}{E_{1}}\right).$
To see (*), note that
@@
With a similar proof, we can restate the following versions of Theorems D.13 and D.15 replacing $\operatorname{{\rm logistic}}$ by $\operatorname{{\rm svm}}$.
Theorem D.22.
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E_{1}$,
@@
D.4 Coreset for $\operatorname{{\rm ReLU}}$ Function
@@
For $\operatorname{{\rm ReLU}}(t)=\max(0,t)$, we derive our results from those in previous section using a simple trick. Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E^{2}_{1}$ or $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$. It is easy to verify that $\mathsf{K}^{\prime}(\cdot)=1+\mathsf{K}(\cdot)$ is also a reproducing kernel associated with $H^{\prime}$ with $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H^{\prime}}^{2}\leq(1+E_{1})^{2}$ or $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H^{\prime}}\geq 1+A\right\}\right)=0$. As $\operatorname{{\rm svm}}(1+t)=\max(0,-t)$, if we plug $\mathsf{K}^{\prime}$ in the results given in the previous section, we can replicate them for $\phi(t)=\max(0,-t)$ by replacing $E_{1}$ and $A$ with $(1+E_{1})$ and $(1+A)$. While $\phi$ is not precisely the $\operatorname{{\rm ReLU}}$ function, for the case that the kernel is the standard linear kernel in Euclidean space,
$\ell_{\phi,k}(x,w)=\max(0,-\langle x,w\rangle)+\frac{1}{k}\|w\|_{2}^{2}=\max(0,\langle x,-w\rangle)+\frac{1}{k}\|w\|_{2}^{2}=\ell_{\operatorname{{\rm ReLU}},k}(x,-w).$